Calculate the mass of the earth if the radius is approximately 6400km, if g=9.8m/s² and G=6.7×10–11Nm²kg–²


1
Expert's answer
2021-01-07T10:55:11-0500
"g=\frac{GM}{R^2}\\9.8=\frac{(6.7\cdot10^{-11})M}{(6400000)^2}\\M=6\cdot10^{24}\ kg"